Hello, I am working with functions just substituting and solving problems. The given is f(x)=5x^2+4x-3 and what I am trying to solve is f(x+1) put into the equation. I understand that to substitute the problem would be f(x+1)=5(x+1)^2+4(x+1)-3 I am just unsure of how to solve. If you could help that would be great.
Found 2 solutions by josgarithmetic, jim_thompson5910:
Answer by josgarithmetic(39620)

Best suggestion, do the indicated multiplications and then any simplifications which become possible. Put the result into general form of terms of x of decreasing degree.

5(x+1)^2 becomes what?
4(x+1) becomes what?

... and then "-3".
...

Answer by jim_thompson5910(35256)

f(x)=5x^2+4x-3

f(x+1)=5(x+1)^2+4(x+1)-3 ... so far, so good

Now you expand, distribute, and combine like terms to simplify

f(x+1)=5(x+1)(x+1)+4(x+1)-3

f(x+1)=5(x^2+2x+1)+4(x+1)-3

f(x+1)=5x^2+10x+5+4x+4-3

f(x+1)=5x^2+14x+6


So the answer is f(x+1)=5x^2+14x+6